Find the probability of at least 2girls in 9 births. Assume male/female are equally likely and births are independent events.
Answer by stanbon(75887)

Find the probability of at least 2girls in 9 births. Assume male/female are equally likely and births are independent events.
--------------
Binomial Problem with n = 9 and p(girl) = 1/2
--------
P(2<= x <=9) = 1 - P(0<= x <=1) = 1 - binomcdf(9,1/2,1) = 0.9805
===============
Cheers,
Stan H.
===============